"Suppose that the length of a radius of a circle is the same as the length of a side of a square. If the circumference of the circle is 11.4 centimeters longer than the perimeter of the square, find the length of a radius of the circle. (Use 3.14 as an approximation for "Pie")."
Can I get step by step help with this problem? Not sure how to set up the equation or even solve.
Thank you!

Answer by mananth(16946)

circle radius = r
circumference = 2*pi*r
length of side of square= r
perimeter = 4r
circumference- perimeter = 11.4

6.28r-4r=11.4
2.28r=11.4
r= 11.4/2.28
= 5 cm
radius = 5 cm